ย A hardware manufacturer produces bolts used to assemble various machines. Assume that the diameter of bolts produced by this manufacturer has an unknown population mean ๐ and the standard deviation is 0.1 mm. suppose the average diameter of a simple random sample of 50 bolts is 5.11 mm. (a) Calculate the margin of error of a 95% confidence interval for ๐.


1
Expert's answer
2022-05-24T09:37:05-0400

The critical value forย "\alpha = 0.05" isย "z_c = z_{1-\alpha/2} = 1.96."


"ME=z_c\times\dfrac{\sigma}{\sqrt{n}}=1.96\times\dfrac{0.1}{\sqrt{50}}""=0.0277"